# Improving Sensitivity of Vector-like Top Partner Searches with Jet Substructure

### v1

@@ I Introduction
The search for new physics beyond the standard model (BSM) remains one of the most important goals of current and future high-energy collider experiments. Despite many successes, the standard model (SM) of elementary particles leaves several fundamental questions unanswered, such as the nature of electroweak symmetry breaking Giudice (2013); Alekhin et al. (2012); Degrassi et al. (2012), the origin of mass hierarchies Froggatt and Nielsen (1979); Fritzsch and Xing (2000), matter-antimatter asymmetry Riotto and Trodden (1999); Dine and Kusenko (2003), the composition of dark matter Clowe et al. (2006); Sofue and Rubin (2001), etc. Many BSM models Perelstein et al. (2004); Matsedonskyi et al. (2013) propose the existence of new resonances, such as vector-like fermions Aguilar-Saavedra et al. (2013); Falkowski et al. (2014); del Aguila et al. (1990), that can provide critical insights into some of these open problems and offer experimentally verifiable predictions at the TeV scale Hayrapetyan et al. (2025). Vector-like fermions are directly related to the stability of the electroweak vacuum Hiller et al. (2022). A vector-like top partner, being a color triplet and having the same electric charge as the SM top quark, couples to the Higgs boson and alters the renormalization group evolution of the Higgs quartic coupling. Its Yukawa interaction contributes positively to the running of the Higgs quartic coupling, while mixing with the SM top reduces the top Yukawa contribution that typically drives the coupling negative. Together, these effects slow down or prevent the Higgs quartic coupling from turning negative at high energy scales, thus stabilizing the electroweak vacuum all the way up to the Planck scale. In addition, colored vector-like fermions can mediate interactions between dark matter and SM quarks as indicated in the phenomenological analyses of vector-like portal dark matter models Colucci et al. (2018); Ghosh and Konar (2024); Ghosh et al. (2022); Ghosh and Konar (2025). Thus identifying such states would not only signify the discovery of new physics phenomena; but also provide a profound understanding of the structure of fundamental interactions. Extraction of rare BSM signals from overwhelming SM backgrounds requires improved analysis techniques and theoretical modelling. State-of-the-art jet substructure methods Marzani et al. (2019); Kogler et al. (2019), clustering algorithms Salam (2010), and machine learning techniques Duarte (2025); Radovic et al. (2018) have become indispensable tools for identifying boosted heavy resonances and complex event topologies that are characteristic of various BSM signatures at the Large Hadron Collider (LHC). At the same time, access to high-quality data samples, such as CMS Open Data CMS ; CER (2020), allows for a more reliable background model and realistic performance estimates.
In this work, we investigate the production of a heavy vector-like top partner ($T$) in association with a SM top quark, mediated through chromomagnetic interactions Buchkremer et al. (2013); Choudhury et al. (2021). The top partner decays through four main channels: $tZ$, $th$, $tg$, and $bW$; each leading to distinct final state signatures. Among these, direct searches at the LHC have been performed primarily in the $tZ$ and $th$ channels, leading to stringent lower limits on the mass of the top partner around a few TeV Aad et al. (2024); Tumasyan et al. (2023a); Hayrapetyan et al. (2024). Furthermore, indirect constraints on the BSM nature of the $tg$ coupling have been derived using $\mathrm{t\bar{t}}$ events Hayreter and Valencia (2016); Tumasyan et al. (2023b). However, only a handful of direct searches in the $bW$ decay mode have been performed at the LHC with full Run-II data Tumasyan et al. (2023c) due to its limited sensitivity. Hence, we put careful focus on the $bW$ mode of the $T$ quark decay along with the usual modes in this article.
@@
The nature of modern BSM searches require exploration at increasingly higher energy regimes. At these high energies, fat jets frequently arise from the hadronic decays of boosted heavy particles such as top quarks or heavy bosons, after parton showering and hadronization. One of the primary aspect of this paper is to assess the efficacy of the recently proposed “Dynamic radius” (DR) jet clustering algorithm Mukhopadhyaya et al. (2023a), which is potentially advantageous in the higher energy regime where such fat jets are prominent. The substructure of the fat jets originating from massive boosted particle decays differs significantly from that of narrow QCD jets owing to their decay history,. The DR algorithm leverages this substructure – more precisely, combinations of energy correlation functions Larkoski et al. (2013) to adaptively set the jet radius based on local kinematics, thus producing variable-sized jets in an event. Some recent studies have made use of the DR clustering algorithm with zero pile up sceanrio as indicated in Refs. Kar et al. (2024); Mukhopadhyaya et al. (2023b); Ghosh et al. (2025) However, understanding the performance of this algorithm in realistic experimental conditions is extremely crucial. This is the first study incorporating full detector simulation including a realistic model for pile up, to evaluate the robustness of the DR jet clustering method under practical LHC-like conditions.
@@
The final state for this study consists of exactly one lepton, exactly one fat jet, at least one narrow jet that is $b~{}$-tagged, and missing transverse momentum ($\text{p}_{\text{T}}^{\text{miss}}$), capturing event topologies sensitive to the presence of vector-like top partners. The probe for the signal process is conducted using jets clustered using traditional fixed-radius as well as the more recent dynamic radius approaches. We employ powerful jet substructure tools, namely Soft Drop (SD) grooming Larkoski et al. (2014a), $N~{}$-subjettiness Thaler and Van Tilburg (2011), and energy correlation functions (ECFs) Larkoski et al. (2013) to better understand the nature and the behavior of the fat jet in the event. In addition, we also studied several event shape variables, such as Sphericity Bjorken and Brodsky (1970); Donoghue et al. (1979), Aplanarity Aad et al. (2012), Fox-Wolfram moments ($H$Fox and Wolfram (1979); Bernaciak et al. (2013) etc., along with lepton and jet kinematic information in a multivariate analysis (MVA), to effectively distinguish signal events from the background. The signal events are simulated with MadGraph5_aMC@NLO Alwall et al. (2014); Frederix et al. (2018) together with Pythia8 Bierlich et al. (2022) for hadronization and parton showering and Delphes de Favereau et al. (2014) for detector response modelling. For SM backgrounds, high statistics Monte Carlo (MC) samples from CMS Open Data have been utilized for a reliable estimate. This study also includes pileup simulation with a realistic value of the mean pileup interactions, considering the Run-II LHC conditions.
@@
The article is organized as follows. Section II.1 introduces the signal model based on chromomagnetic coupling between the vector-like top partner and the SM top quark, while the description of signal and background simulation procedures are presented in Sections II.2 and II.3. The DR algorithm and relevant jet substructure techniques are presented in Section III. Section IV discusses the event reconstruction and selection strategy. In Section V, we detail the multivariate analysis strategy, optimization methods, and present our main results. Finally, Section VI provides a summary of the work and an outlook towards the future prospects.
@@ II.1 Signal model
The vector-like $T$ quark is a triplet under $SU(3)$, and its electric charge is determined by its hypercharge and isospin structure. Additionally, it can couple with any generation of the SM quark. We assume the heavy top partner predominantly interacts with the third-generation SM quarks. The $T$ quark is a singlet under weak isospin and has the same electromagnetic charge as the top quark; hence, it mixes with the top quark. The general Lagrangian terms are given below Buchkremer et al. (2013); Choudhury et al. (2021).
@@
Here, $t$ and $h$ denote the SM top quark and higgs boson, respectively. $P_{L/R}=\dfrac{1\mp\gamma_{5}}{2}$ represent the left and right-handed projection operators. The mixing angle between the left-handed (right-handed) SM top quark and the left-handed (right-handed) $T$ quark is denoted by $\theta_{L}$ ($\theta_{R}$). The symbols $M_{T}$, $m_{t}$, and $M_{W}$ represent the masses of the $T$ quark, the SM top quark, and the $W$ boson, respectively. The $SU(2)_{L}$ gauge coupling is $g_{w}$; $s_{w}$ and $c_{w}$ are sine and cosine of the electroweak mixing angle. The single production of the exotic $T$ quark in association with the SM top quark, is possible at the LHC via gluon fusion. The effective Lagrangian of such interactions is as follows.
@@
where $\sigma_{\mu\nu}=i/2~{}(\gamma_{\mu}\gamma_{\nu}-\gamma_{\nu}\gamma_{\mu})$, and $g_{s}$ is the SM strong coupling constant. The notation $G^{a}_{\mu\nu}$ represents the gluon field strength tensor, and $\tau^{a}$ corresponds to the $SU(3)$ generators. $C_{tRL}$ and $C_{tLR}$ are the coupling strengths. We set the new physics scale $\Lambda=8$ TeV for our phenomenological analysis. A smaller or larger value of $\Lambda$ will result in a correspondingly larger or smaller production cross-section, respectively. However, the signal efficiency remains unchanged as long as the mass of the $T$ quark is fixed.
Figure 1: Variation of cross section with the coupling (left) and with the top-partner ($T$) mass (middle) for the associated production of $T$ with the SM top quark at the 13 TeV LHC are presented. In these plots, we set the new physics scale $\Lambda=8~{}$ TeV. For the left plot, $M_{T}$ is set at 2.22.22.2~{}2.2TeV, whereas the couplings $C_{tRL}=C_{tLR}$ are set to $0.8$ for the middle plot. The variation of the branching ratio with the mixing angle is presented for all four decay modes of the top-partner in the right. Here we set $M_{T}=2.2~{}$TeV, $C_{tRL}=C_{tLR}=0.8$, and $\Lambda=8~{}$ TeV.
Fig. 1 shows the variation of the partonic cross section for $T$-quark production in association with a top quark at the 13 TeV LHC, along with the branching ratios of its various decay channels, for a fixed new physics scale of $\Lambda=8$ TeV. As seen in Eqs.(1)–(3), when the mixing angles are very small ($\theta_{L},\theta_{R}\to 0$), the branching ratios for $T\to tZ$, $T\to bW$, and $T\to th$ become negligibly small, and the $T$ quark decays predominantly into a top quark and a gluon. However, with increasing mixing between the $T$ quark and the top quark, the branching fractions into the electroweak channels $T\to tZ$, $T\to bW$, and $T\to th$ increase significantly.
@@ II.2 Signal simulation
The process we are interested in is the production of a SM top quark and its vector-like partner $T$ via the effective chromomagnetic interaction.
$pp$$\to$ $t\,\overline{T}+X,\quad{\rm and}$ (5)
$pp$$\to$ $T\,\overline{t}+X.$ (6)
@@
A set of representative parton-level Feynman diagrams corresponding to the process has been given in Fig. 2.
@@
Figure 2: Representative parton level Feynman diagrams for the process $pp\to T\bar{t}+\bar{T}t$.
For further study and discussion of our result, we chose a benchmark scenario as follows:
$\frac{C_{tRL}}{\Lambda}=\frac{C_{tLR}}{\Lambda}=0.1~{}\text{TeV}^{-1},\qquad\theta_{L}=\theta_{R}=0.06~{}\text{radian},\qquad\text{and}\ M_{T}=2.2~{}{\rm TeV}.$ (7)
@@
To simulate the processes indicated by Eqs. (5) and (6), we first implemented all the interaction terms defined in Eqs.(1)–(4), along with the SM Lagrangian, using FeynRules Christensen and Duhr (2009); Alloul et al. (2014) to generate a Universal FeynRules Output (UFO) Degrande et al. (2012) model file. This UFO model was then used in the MadGraph5_aMC@NLO framework to generate parton-level events at a centre-of-mass energy of 13 TeV with the parton distribution function (PDF) set to NNPDF30_nlo_nf_5_pdfas PDFset Ball et al. (2015). The factorization and renormalization scales are set to the default dynamical scale in MadGraph5, which corresponds to the transverse mass of the $2\to 2$ scattering process at leading order (LO). The generated events are then passed through Pythia8 Bierlich et al. (2022) for parton showering and hadronization. In the Pythia8 simulation, the CMS Tune MonashStar Khachatryan et al. (2016), based on Monash 2013 tune Skands et al. (2014), has been used to better match CMS data, especially to model the underlying event, initial-state radiation, and beam remnants. Finally, a fast detector simulation is performed using Delphes de Favereau et al. (2014) to incorporate detector effects. The parton-level kinematics of the produced top partner is shown in Fig. 3 for various mass hypotheses. The higher tail of the $\text{p}_{\text{T}}~{}$–spectrum of the $T$ quark has a distinct dependence on its mass. The cross section of the signal process, the $T$ quark decay width, and its branching fractions for all of the four decay modes are is listed in Appendix B for different mass hypotheses.
@@
Figure 3: Parton-level distributions of $\text{p}_{\text{T}}$ (left) and $\eta$ (right) of the top partner for different mass hypotheses.
@@
Depending on the decay mode of the top partner into $tZ$, $th$, $tg$, or $bW$, the above processes could be realized into different event categories. To suppress a huge QCD multijet background at the LHC, we focus on the semi-leptonic channels requiring one lepton ($\ell\ =\ e,\mu$) in the final state. Therefore, the final states where one of the $W$’s originating from either of the $t$, $\bar{t}$, $T$, or $\bar{T}$ decay leptonically, are studied. The event sample for our search can be broadly classified into following two categories:
${\rm Two\ fat\ jets\ category:}$ $1\ell+2\ {\rm fat\ jets}+b{\rm-jet}+\text{p}_{\text{T}}^{\text{miss}},$ (8)
${\rm One\ fat\ jet\ category:}$ $1\ell+1\ {\rm fat\ jet}+b{\rm-jet(s)}+\text{p}_{\text{T}}^{\text{miss}}.$ (9)
@@
Table 1 lists the different decay modes of the top partner and their possible contribution to each event category. The signal fractions are larger in the two fat jets event category for all decay modes. However, in this event category, the $\mathrm{t\bar{t}}$ + X (X = W, Z, jets etc.) backgrounds also have significantly larger contribution leading to poor signal-to-background ratio and thus making it unsuitable as a search region for the signal process. Therefore, we focus on the one fat jet category where the contribution of the $bW$ decay mode is relatively higher.

### v2

@@ I Introduction
The search for new physics beyond the standard model (BSM) remains one of the most important goals of current and future high-energy collider experiments. Despite many successes, the standard model (SM) of elementary particles leaves several fundamental questions unanswered, such as the nature of electroweak symmetry breaking and the stability of the electroweak vacuum Giudice:2013yca ; Arkani-Hamed:2012dcq ; Alekhin:2012py ; Degrassi:2012ry ; Dawson:2010jx , the origin of mass hierarchies Froggatt:1978nt ; Fritzsch:1999ee , matter-antimatter asymmetry Riotto:1999yt ; Dine:2003ax , the composition of dark matter Clowe:2006eq ; Sofue:2000jx , etc. Many BSM models Perelstein:2003wd ; Matsedonskyi:2012ym ; Chanowitz:1978mv propose the existence of new resonances, such as vector-like fermions Aguilar-Saavedra:2013qpa ; Falkowski:2013jya ; Dawson:2012di ; delAguila:1989rq , that can provide critical insights into some of these open problems and offer experimentally verifiable predictions at the TeV scale CMS:2024bni . Vector-like fermions are directly related to the stability of the electroweak vacuum Hiller:2022rla . A vector-like top partner, being a colour triplet and having the same electric charge as the SM top quark, couples to the Higgs boson and alters the renormalization group evolution of the Higgs quartic coupling. Its Yukawa interaction contributes positively to the running of the Higgs quartic coupling, while its mixing with the SM top quark reduces the Yukawa contribution due to the top quark that typically drives the coupling negative. Together, these effects slow down or prevent the Higgs quartic coupling from turning negative at high energy scales, thus stabilizing the electroweak vacuum all the way up to the Planck scale Hiller:2022rla . In addition, coloured vector-like fermions can mediate interactions between dark matter and SM quarks as indicated in the phenomenological analyses of vector-like portal dark matter models Ghosh:2024boo ; Colucci:2018vxz ; Ghosh:2025agw ; Ghosh:2022rta ; Ghosh:2023xhs . Thus, identifying such BSM states would not only signify the discovery of new physics phenomena but also provide a profound understanding of the structure of fundamental interactions.
Extraction of rare BSM signals from overwhelming SM backgrounds requires improved analysis techniques and theoretical modelling. State-of-the-art jet substructure methods Larkoski:2017jix ; Marzani:2019hun ; Kogler:2018hem , clustering algorithms Salam:2010nqg ; Sapeta:2015gee , and machine learning techniques Duarte:2024lsg ; Radovic:2018dip have become indispensable tools for identifying boosted heavy resonances and complex event topologies that are characteristic of various BSM signatures at the Large Hadron Collider (LHC). At the same time, access to high-quality data and simulation samples, such as CMS Open Data CMS_Open_Data ; CERN-OPEN-2020-013 , allows for a more reliable background model and realistic performance estimates.
@@
The nature of modern BSM searches requires exploration at increasingly higher energy regimes. At these high energies, fat jets frequently arise from the hadronic decays of boosted heavy particles such as the top quarks or the heavy bosons, after parton showering and hadronization. One of the primary aspects of this paper is to assess the efficacy of the recently proposed “Dynamic radius” (DR) jet clustering algorithm Mukhopadhyaya:2023rsb , which is advantageous in the higher energy regime where such fat jets are prominent. The substructure of the fat jets originating from massive boosted particle decays differs significantly from that of narrow QCD jets Sterman:1977wj owing to their decay history. The DR algorithm leverages this substructure – more precisely, combinations of energy correlation functions Larkoski:2013eya to adaptively set the jet radius based on local kinematics, thus producing variable-sized jets in an event. Some recent studies have made use of the DR clustering algorithm with zero pile up scenario Kar:2022hxn ; Mukhopadhyaya:2023akv ; Ghosh:2025gue . However, understanding the performance of this algorithm in realistic experimental conditions remains extremely crucial.
@@
In this work, we investigate the production of a heavy vector-like top partner ($T$) in association with a SM top quark, mediated through chromomagnetic interactions Buchkremer:2013bha ; Choudhury:2021nib ; Balaji:2021lpr . The top partner decays through four main channels: $\rm tZ$, $\rm th$, $\rm tg$, and $\rm bW$; each leading to distinct final state signatures. [^1] In this article, we followed a convention to denote all SM particles in roman and BSM particles in italics. However, the associated fields of both SM and BSM particles are denoted in italics. Among these, direct searches at the LHC have been performed primarily in the $\rm tZ$ and $\rm th$ channels, leading to stringent lower bounds on the mass of the top partner around a few TeV ATLAS:2024xne ; CMS:2023agg ; CMS:2024qdd . Furthermore, indirect constraints on the BSM nature of the $\rm tg$ coupling have been derived using $\mathrm{t\bar{t}}$ events Hayreter:2015ryk ; CMS:2022voq . However, only a handful of direct searches in the $\rm bW$ decay mode have been performed at the LHC with full Run-II data CMS:2022fck ; CMS-PAS-B2G-23-009 due to its limited sensitivity. The final state for this study consists of exactly one lepton, exactly one fat jet, at least one narrow jet that is b-tagged, and missing transverse momentum ($\text{p}_{\text{T}}^{\text{miss}}$), capturing event topologies sensitive to the presence of vector-like top partners. The bW decay mode of the $T$ quark constitutes a significant contribution to this specific final state. The probe for the signal process is conducted using jets clustered using traditional fixed-radius as well as the more recent DR approaches. We employ powerful jet substructure tools, namely Soft Drop (SD) grooming Larkoski:2014wba , $N$-subjettiness Thaler:2010tr , and energy correlation functions (ECFs) Larkoski:2013eya to better understand the nature and the behaviour of the fat jet in the event. In addition, we also studied several event shape variables, such as Sphericity Bjorken:1969wi ; Donoghue:1979vi , Aplanarity ATLAS:2012tch , Fox-Wolfram moments (HHFox:1978vw ; Bernaciak:2012nh etc., along with lepton and jet kinematic information in a multivariate analysis (MVA), to effectively distinguish signal events from the background. The signal events are simulated with MadGraph5_aMC@NLO Alwall:2014hca ; Frederix:2018nkq together with Pythia8 Bierlich:2022pfr for hadronization and parton showering and Delphes deFavereau:2013fsa for detector response modelling. For SM backgrounds, high statistics Monte Carlo (MC) samples from CMS Open Data have been utilized for a reliable estimate. This study also includes pile up simulation with a realistic value of the mean pileup interactions, considering the Run-II LHC conditions.
Before proceeding further, it should be noted that the main goal of this work is to explore the usefulness of the DR jet clustering in highly boosted final states and to compare its performance with standard fixed-radius clustering. Impact of various uncertainty sources, such as pile up contamination removal, jet energy scale and resolution, hadronization model etc., on the DR algorithm require dedicated in-depth studies that is beyond the scope of the present work. This study is intended as a proof-of-concept, focusing on qualitative improvements achievable with DR jet clustering using a purely statistical approach. This is an initial step to demonstrate the applicability of the DR jet clustering, considering a well-motivated candidate BSM process involving a heavy top partner coupled strongly with the SM top quark via chromomagnetic coupling.
@@
The article is organized as follows. Section II.1 introduces the signal model based on chromomagnetic coupling between the vector-like top partner and the SM top quark, while the description of signal and background simulation procedures is presented in Sections II.2 and II.3, respectively. The DR algorithm and relevant jet substructure techniques are presented in Section III together with a control study using $\mathrm{t\bar{t}}$ + jets events. Section IV discusses the event reconstruction and selection strategy. In Section V, we detail the multivariate analysis strategy, optimization methods, and present our main results. Finally, Section VI provides a summary of the work and an outlook towards the future prospects.
@@ II.1 Signal Model
The vector-like $T$ quark is a triplet under $SU(3)$, and its electric charge is determined by its hypercharge and isospin structure. Additionally, it can couple with any generation of the SM quark. We assume the heavy top partner predominantly interacts with the third-generation SM quarks. The $T$ quark is a singlet under weak isospin and has the same electromagnetic charge as the top quark; hence, it mixes with the top quark. The general Lagrangian terms are given below Buchkremer:2013bha ; Choudhury:2021nib ; Balaji:2021lpr .
@@
Here, $t$ and $h$ denote the SM top quark and higgs boson, respectively. $P_{L/R}=\dfrac{1\mp\gamma_{5}}{2}$ represent the left- and right-chiral projection operators. The mixing angle between the left-chiral (right-chiral) SM top quark and the left-chiral (right-chiral) $T$ quark is denoted by $\theta_{L}$ ($\theta_{R}$). The symbols $M_{T}$, $m_{t}$, and $M_{W}$ represent the masses of the $T$ quark, the SM top quark, and the WW boson, respectively. The $SU(2)_{L}$ gauge coupling is $g_{w}$; $s_{w}$ and $c_{w}$ are sine and cosine of the electroweak mixing angle. The single production of the exotic $T$ quark in association with the SM top quark, is possible at the LHC via gluon fusion. The effective Lagrangian of such interactions is as follows.
@@
where $\sigma_{\mu\nu}=i/2~(\gamma_{\mu}\gamma_{\nu}-\gamma_{\nu}\gamma_{\mu})$, and $g_{s}$ is the SM strong coupling constant. The notation $G^{a}_{\mu\nu}$ represents the gluon field strength tensor, and $\tau^{a}$ corresponds to the $SU(3)$ generators. The notation $\Lambda$ denotes the new physics scale, and $C_{tRL}$, $C_{tLR}$ are the coupling strengths. Corresponding to the above Lagrangian given in Eqs. (1)–(4), the $T$ quark has four decay channels, namely bW, tZ, th, and tg. While the Eq. (4) primarily contributes to the production of the $T$ quark, it is also responsible for the tg decay mode. From Eqs. (1)–(3), one notices that the branching ratios for $T\to\rm tZ$, $T\to\rm bW$, and $T\to\rm th$ become negligibly small when the mixing angles are very small ($\theta_{L},\theta_{R}\to 0$), and the $T$ quark decays predominantly into a top quark and a gluon. However, with increasing mixing between the $T$ quark and the top quark, the branching fractions into the electroweak channels $T\to\rm tZ$, $T\to\rm bW$, and $T\to\rm th$ increase significantly.
@@ II.2 Signal Simulation
The process we are interested in is the associated production of an SM top quark and its vector-like partner $T$ via the effective chromomagnetic interaction.
$\mathrm{pp}$ $\to$ $\mathrm{t}\,\overline{T}+X,\quad{\rm and}$ (5)
$\mathrm{pp}$ $\to$ $T\,\overline{\rm t}+X.$ (6)
@@
A set of representative parton-level Feynman diagrams corresponding to the process has been given in Figure 1.
@@
Figure 1: Representative parton-level Feynman diagrams for the process $\mathrm{pp}\to T\,\bar{\mathrm{t}}\,/\,\bar{T}\,\mathrm{t}$.
@@
To simulate the processes indicated by Eqs. (5) and (6), we first implemented all the interaction terms defined in Eqs. (1)–(4), along with the SM Lagrangian, using FeynRules Christensen:2008py ; Alloul:2013bka to generate a Universal FeynRules Output (UFO) Degrande:2011ua model file. This UFO model was then used in the MadGraph5_aMC@NLO framework to generate parton-level events at a centre-of-mass energy of $\sqrt{s}~=~13$~TeV with the parton distribution function (PDF) set to NNPDF30_nlo_nf_5_pdfas PDFset NNPDF:2014otw . The factorization and renormalization scales are set to the default dynamical scale in MadGraph5, which corresponds to the transverse mass of the $2\to 2$ scattering process at leading order (LO). Figure 2 shows the variation of the partonic cross section for $T$ quark production in association with a top quark at $\sqrt{s}~=~13$~TeV, along with the branching ratios of its various decay channels, for a fixed new physics scale of $\Lambda=8$ TeV.
Figure 2: Cross section vs. coupling (upper left) and vs. $T$ quark mass (upper right) are presented for $\mathrm{pp}\to T~\bar{\mathrm{t}}~/~\bar{T}~\rm t$ process at $\sqrt{s}~=$~13 TeV with new physics scale $\Lambda~=~8$~TeV, $M_{T}~=~2.2$~TeV (upper left), and $C_{tRL}~=~C_{tLR}~=~0.8$ (upper right). Branching ratios (BRs) vs. mixing angle for all $T$ decay modes (lower) are shown with $M_{T}~=~2.2$~TeV, $C_{tRL}~=~C_{tLR}~=~0.8$, and $\Lambda~=~8$~TeV.
To illustrate the collider analysis for this study, we consider a benchmark point that is consistent with existing experimental bounds. The strongest lower bound on $M_{T}$ is about 1.8 TeV ATLAS:2024xne . We further consider $\theta_{L}=\theta_{R}$ and $C_{tRL}=C_{tLR}$ to reduce the number of free parameters. The couplings $C_{tRL}$ and $C_{tLR}$ control the top-partner pair production cross section, which is restricted to $\mathcal{O}(1~{\rm fb})$ CMS:2024bni , resulting in an upper limit on $\frac{C_{tRL}}{\Lambda}=\frac{C_{tLR}}{\Lambda}\lesssim 0.6~\text{TeV}^{-1}$. Furthermore, the mixing parameters $\theta_{L}$ and $\theta_{R}$ affect the Higgs boson branching ratio to gluon-gluon and di-photon modes. One finds $\theta_{L}=\theta_{R}\lesssim 0.25$ radian for $M_{T}=2.2$ TeV from Refs. ATLAS:2025hhd ; Dawson:2012di ; Zhang:2025zkn . Based on these factors, we choose the benchmark scenario as follows:
$\frac{C_{tRL}}{\Lambda}=\frac{C_{tLR}}{\Lambda}=0.1~\text{TeV}^{-1},\qquad\theta_{L}=\theta_{R}=0.06~\text{radian},\qquad\text{and}\quad\ M_{T}=2.2~{\rm TeV}.$ (7)
The parton-level kinematics of the produced top partner are shown in Fig. 3 for different mass hypotheses. The higher tail of the $\text{p}_{\text{T}}$-spectrum of the $T$ quark has a distinct dependence on its mass. The cross section of the signal process, the $T$ quark decay width, and its branching fractions for all of the four decay modes are listed in Appendix A for different mass hypotheses.
@@
Figure 3: Distributions of the top partner $\text{p}_{\text{T}}$ (left) and $\eta$ (right) at parton level for different mass hypotheses for the process $\mathrm{pp}\to T\,\bar{\mathrm{t}}\,/\,\bar{T}\,\mathrm{t}$ at $\sqrt{s}=13$ TeV.
The generated events are then passed through Pythia8 Bierlich:2022pfr for parton showering and hadronization. In the Pythia8 simulation, the CMS Tune MonashStar CMS:2015wcf , based on Monash 2013 tune Skands:2014pea , has been used to better match CMS data, especially to model the underlying event, initial-state radiation, and beam remnants. We then incorporated the effects of pile up in the signal MC samples by overlaying minimum-bias event samples with the hard-scattered events with approximately $20$ average pile up interactions per event. Finally, a fast detector simulation is performed on the pile up overlaid samples using Delphes_v3.5.0 deFavereau:2013fsa that employs an overall detector geometry definition and a parametric model of average detector response [^2] The choice of using Delphes fast detector simulation for signal MC samples was made owing to limitations in the available computing and storage resources at our disposal for this study.. The signal event simulation details are summarized in Table 1.
Table 1: Details of signal simulation tools and associated settings.
BSM model SM + Eqs. (1)–(4) using FeynRules Christensen:2008py ; Alloul:2013bka and
Universal FeynRules Output (UFO) Degrande:2011ua
Matrix element Eqs. (5) and (6) using MadGraph5_aMC@NLO Alwall:2014hca
at LO with $\sqrt{s}=13$ TeV
PDF set NNPDF30_nlo_nf_5_pdfas NNPDF:2014otw
Parton showering & Hadronization Pythia8 Bierlich:2022pfr
Underlying event CMS Tune MonashStar CMS:2015wcf , based on Monash
2013 tune Skands:2014pea
Detector simulation model Delphes_v3.5.0 deFavereau:2013fsa
@@
Depending on the decay mode of the top partner into $\rm tZ$, $\rm th$, $\rm tg$, or $\rm bW$, the above processes could be realized into different event categories. To suppress a huge QCD multijet background at the LHC, we focus on the semi-leptonic channels requiring one lepton ($\ell\ =\ e,\mu$) in the final state. Therefore, the final states where one of the $\rm W$’s originating from either of the $\rm t$, $\rm{\bar{t}}$, $T$, or $\bar{T}$ decay leptonically, are studied. The event sample for our search can be broadly classified into the following two categories:
${\rm Two\ fat\ jets\ category:}$ $1\ell+2\ {\rm fat\ jets}+1\ {\rm b-jet}+\text{p}_{\text{T}}^{\text{miss}},$ (8)
${\rm One\ fat\ jet\ category:}$ $1\ell+1\ {\rm fat\ jet}+\geqslant 1\ {\rm b-jet(s)}+\text{p}_{\text{T}}^{\text{miss}}.$ (9)
@@
Table 2 lists the different decay modes of the top partner and their possible contribution to each event category applying the same selection criteria as defined in Section IV on the lepton and jets. The signal fractions are larger in the two fat jets event category for all decay modes. However, in this event category, the $\mathrm{t\bar{t}}$ + X (X = W, Z, jets, etc.) backgrounds also have a significantly larger contribution, leading to a poor signal-to-background ratio and thus making it unsuitable as a search region for the signal process. Therefore, we focus on the one fat jet category where the contribution of the $\rm bW$ decay mode is relatively higher.